# Coupled Input-Output Dimension Reduction: Application to goal-oriented Bayesian experimental design and global sensitivity analysis

### v1

@@ Abstract
We introduce a new method to jointly reduce the dimension of the input and output space of a high-dimensional function. Choosing a reduced input subspace influences which output subspace is relevant and vice versa. Conventional methods focus on reducing either the input or output space, even though both are often reduced simultaneously in practice. Our coupled approach naturally supports goal-oriented dimension reduction, where either an input or output quantity of interest is prescribed. We consider, in particular, goal-oriented sensor placement and goal-oriented sensitivity analysis, which can be viewed as dimension reduction where the most important output or, respectively, input components are chosen. Both applications present difficult combinatorial optimization problems with expensive objectives such as the expected information gain and Sobol’ indices. By optimizing gradient-based bounds, we can determine the most informative sensors and most sensitive parameters as the largest diagonal entries of some diagnostic matrices, thus bypassing the combinatorial optimization and objective evaluation.
@@ 2 Problem Setting and Objectives
with $\operatorname{\mathbb{D}_{KL}}(\nu\|\mu)=\int\log(\nu/\mu)\mathop{}\!\mathrm{d}\nu$. For (O2) to be close to 00, the parameter $X$ needs to be essentially independent of the observations in the subspace $\operatorname{Ker}(V_{s})$, and observations of $Y$ should primarily provide information within the parameter subspace $\operatorname{Im}(U_{r})$. To sample from the approximate density $\widetilde{\pi}_{X|Y}$, we can independently sample $X_{r}\in\mathbb{R}^{r}$, which depends on the projected observation (see algorithms to sample the reduced posterior in [12, Section 6]), and its orthogonal complement $X_{\perp}\in\mathbb{R}^{d-r}$, which only depends on the prior.
Bayesian optimal experimental design (BOED) aims to select a set of sensors $\tau\subseteq\{1,\ldots,m\}$ with $|\tau|=s$ that maximizes the information gained about the parameter that is to be inferred. Measurements from the sensors $\tau$ are gathered in the vector $Y_{\tau}=V_{\tau}^{\top}Y$, where $V_{\tau}=[e_{\tau}^{m}]\in\mathbb{R}^{m\times s}$ contains the canonical basis vectors $e_{i}^{m}\in\mathbb{R}^{m}$ indexed by $\tau$. The information content of the experimental design $V_{\tau}$ can be quantified by the expected information gain (EIG), which is defined as the KL-divergence between posterior and prior in expectation over the data $\Phi(V_{\tau})=\mathbb{E}_{Y}[\operatorname{\mathbb{D}_{KL}}(\pi_{X|Y_{\tau}}\|\,\pi_{X})]$. When the inference problem focuses on a specific parameter of interest $X_{r}=U_{r}^{\top}X$, the sensor placement can be improved by maximizing the goal-oriented EIG
(O3) $\max_{V_{\tau}}\;\Phi(V_{\tau}|U_{r})=\mathbb{E}_{Y}\left[\operatorname{\mathbb{D}_{KL}}\left(\pi_{X_{r}|Y_{\tau}}\|\,\pi_{X_{r}}\right)\right],$
@@
Global sensitivity analysis (GSA) aims to identify which set $\tau\subseteq\{1,\ldots,d\}$ of $r=|\tau|$ input variables causes the least output variations in order to fix them to a nominal value and simplify the model. A widely used sensitivity index is the total Sobol’ index $S_{\tau}^{\textup{tot}}=1-\frac{\operatorname{Tr}(\operatorname{Cov}(G(X)|X_{-\tau}))}{\operatorname{Tr}(\operatorname{Cov}(G(X)))}$ with $X_{-\tau}={U_{-\tau}}^{\top}X\in\mathbb{R}^{d-r}$ where $U_{-\tau}=[e_{-\tau}^{d}]\in\mathbb{R}^{d\times(d-r)}$ contains the canonical basis vectors $e_{i}^{d}\in\mathbb{R}^{d}$ indexed by the complement of $\tau$ [24]. A smaller $S_{\tau}^{\textup{tot}}$ implies that the output is less sensitive to the parameters indexed by $\tau$. When there are particular output directions of interest $V_{s}^{\top}G(X)$, it is more suitable to consider minimizing the goal-oriented total Sobol’ index
(O4) $\min_{U_{\tau}}\;S^{\textup{tot}}(U_{\tau}|V_{s})=1-\frac{\operatorname{Tr}\left(\operatorname{Cov}\left(\left.V_{s}^{\top}G(X)\right|U_{-\tau}^{\top}X\right)\right)}{\operatorname{Tr}\left(\operatorname{Cov}\left(V_{s}^{\top}G(X)\right)\right)},$
@@ 3 Coupled Input-Output Dimension Reduction
Proposition 3.3 allows us to obtain optimal $U_{r}$ and $V_{s}$ by minimizing the $L^{2}$-error of the forward model as given by (5), the minimum value for the objective in (O1). After that, one can approximate the low-dimensional function $g^{*}$ using state-of-the-art approximation methods and compute $G^{*}_{\perp}$ to obtain a reduced order model of $G$. We focus on reducing the inputs and/or outputs by minimizing (5) over $U_{r}$ and/or $V_{s}$, which is equivalent to
@@ 3.1 Derivative-based Bounds
We will show in the next subsection that maximizing over either $U_{r}$ or $V_{s}$ yields closed-form solutions that can be computed as dominant eigenvectors of certain diagnostic matrices. There is, however, no closed-form solution for the joint maximization over $U_{r}$ and $V_{s}$. Nonetheless, since the Stiefel manifold $\text{St}(p,q)=\{A\in\mathbb{R}^{p\times q}\;|\;A^{\top}A=I_{q}\}$ is compact [1] and the objective function is continuous with respect to $U_{r}$ and $V_{s}$, the joint maximization problem has a solution. We propose a natural heuristic for its computation in the next subsection. After computing $U_{r}$ and/or $V_{s}$, the lower-bound (11) can be evaluated with no additional cost and allows us to certify the error made by optimizing the upper bound.
The following proposition gives sufficient condition on the density $\pi_{X}$ which ensures $\overline{c}(X)>0$ and $\overline{\mathcal{C}}(X)<\infty$.
@@
While assumption (13) is a classical strict log-concavity assumption on $\pi_{X}$, the assumption (14) means that the density $\pi_{X}$ has to lie above some scaled Gaussian density (see Remark D.5).
@@ 3.2 Minimizing the Error Bound
Computing these matrix-vector products requires only evaluating the linear-tangent operator $u\mapsto\nabla G(X^{(i)})u$ and its adjoint $v\mapsto\nabla G(X^{(i)})^{\top}v$, which are readily available for many applications or can be implemented efficiently, e.g., using finite difference and automatic differentiation.
@@
In a recent paper [12], the authors proposed a joint dimension reduction method for the parameter and data space based on similar diagnostic matrices. Using an information theoretic approach, they derive an upper bound for the posterior error $\mathbb{E}_{Y}[\operatorname{\mathbb{D}_{KL}}(\pi_{X|Y}\|\widetilde{\pi}_{X|Y})]$ which is minimised by $U_{r}^{*}$ and $V_{s}^{*}$ consisting of the dominant eigenvectors of
@@ 4 Application to Goal-oriented Coordinate Selection
4.1 Goal-Oriented Bayesian Optimal Experimental Design (BOED)
@@
Estimating the EIG poses an even greater challenge than solving the Bayesian inference problem as it requires solving the inverse problem itself as a sub-problem. The EIG consists of an expectation over both the posterior and data distributions, necessitating an expensive nested Monte Carlo estimator [47, 14]. Efficient EIG estimation is an active research area, and various approaches exist. Some rely on constructing tractable approximations to the posterior distribution using measure transport [32, 11] or variational inference [22]. Others employ different types of surrogate models [30, 52, 7]. A widely used strategy is to resort to Gaussian linear models for which a closed-form expression of the EIG exists [3, 51, 9]. In addition to the challenging EIG estimation, identifying the optimal set of sensor placements requires solving a combinatorial optimization problem with $\binom{m}{s}$ possible combinations. This NP-hard problem is intractable even for moderate dimensions. Hence, sub-optimal solutions are commonly identified using greedy methods [34, 16, 51] or by solving continuous relaxations of the problem with sparsity constraints [26, 9].
We consider, in particular, goal-oriented BOEDs, where the final target is to infer the parameter in a lower-dimensional subspace of interest. For example, if the PDE model describes a spatio-temporal field with $X$ as the initial state, the initial state in a localized region of the domain could be a parameter of interest. Tailoring the experimental design to maximize the information gained with respect to the parameter of interest allows for more efficient and accurate inference from experimental data. For a given parameter of interest $X_{r}=U_{r}^{\top}X$, we seek the optimal design $V_{\star}$ that maximizes the goal-oriented EIG (see (O3))
@@
(24) $\max_{\begin{subarray}{c}V_{\tau}=[e_{\tau}]\in\mathbb{R}^{m\times s}\\ \tau\subset\{1,\ldots,m\},\;|\tau|=s\end{subarray}}\;\operatorname{Tr}\left(V_{\tau}^{\top}H_{Y}(U_{r})V_{\tau}\right)=\sum_{i\in\tau^{*}}H_{Y}(U_{r})_{ii},$
where the solution $\tau^{*}\subseteq\{1,\ldots,m\}$ is the index set of the $s$ largest diagonal entries of $H_{Y}(U_{r})$. Not only do we bypass computing the EIG, but our solution $\tau^{*}$ can be computed without combinatorial optimization.
@@
The EIM takes as input a set of basis vectors $v_{1},\ldots,v_{s}\in\mathbb{R}^{m}$ and returns a set of indices $\{i_{1},\ldots,i_{s}\}\subseteq\{1,\ldots,m\}$ corresponding to the empirically-best interpolation points for the given basis. Combining the EIM with dimension reduction methods for sensor placement has been considered, for example, in [25, 8]. By using the EIM on the columns of $V_{s}^{*}(U_{r})$, we can thus associate sensor locations $\widetilde{\tau}=\{i_{1},\ldots,i_{s}\}$ to the orthogonal design matrix $V_{s}^{*}(U_{r})$. Compared to $\tau^{*}$, the EIM-based sensors $\widetilde{\tau}$ are less optimal with respect to the goal-oriented EIG lower bound (24). However, our numerical experiments show that the EIM-based sensors often slightly outperform $\tau^{*}$ in terms of the true goal-oriented EIG rather than the lower bound. A reason could be that the sensors indexed by $\widetilde{\tau}$ tend to be less localized than $\tau^{*}$. In the numerical experiments, we also apply the EIM on PCA bases to obtain sensor placements $\widetilde{\tau}^{\textup{PCA}}$, which we will refer to as PCA-EIM sensors.
4.2 Goal-Oriented Sensitivity Analysis (GSA)
Global sensitivity analysis aims to assign an importance value to each set of input variables, reflecting its contribution to the output variations. It is often used for factor prioritization and/or fixing in complex models. Factor prioritization identifies the subset $\tau\subseteq\{1,\ldots,d\}$, $|\tau|=r$ of input parameters causing the most output variation so that the parameters $X_{\tau}=(X_{i})_{i\in\tau}$ can be prioritized e.g., in uncertainty reduction. Factor fixing, on the other hand, identifies the input set $\tau$ with the least influence on the output. Thus, $X_{\tau}$ can be fixed at a nominal value to reduce the complexity of, e.g., parameter estimation problems. For a scalar-valued square-integrable function $f\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$, commonly used sensitivity indices are the closed Sobol’ indices and the total Sobol’ indices defined as
@@ 4.2 Goal-Oriented Sensitivity Analysis (GSA)
respectively, where $X_{-\tau}=(X_{i})_{i\notin\tau}$ and the input variables $X_{i}$ for $i=1,\ldots,d$ are assumed to be independent. The closed Sobol’ index quantifies the portion of output variance attributed to $X_{\tau}$. A higher value implies a stronger influence of $X_{\tau}$ on the output. The total Sobol’ index encompasses output variation due to $X_{\tau}$ along with variations caused by interactions of $X_{\tau}$ with any of the remaining input variables. As a result, it holds that $S^{\textup{cl}}_{\tau}\leq S^{\textup{tot}}_{\tau}$ and $1=S^{\textup{tot}}_{\tau}+S^{\textup{cl}}_{-\tau}$. A lower value in total Sobol’ index suggests that the parameters indexed by $\tau$ have little direct nor indirect impact on the output.
@@
For $V_{s}\in\mathbb{R}^{m\times 1}$ being a vector, the provided indices correspond to the ones for a scalar-valued output (26). Using the Poincaré and Cramér-Rao inequality, we now derive derivative-based bounds for the goal-oriented Sobol’ indices involving similar derivative-based quantities as in Theorem 3.5.
@@
With the above notations, for any matrices $U_{r}\in\mathbb{R}^{d\times r}$ and $V_{s}\in\mathbb{R}^{m\times s}$ with orthogonal columns, we have
(28) $1-\overline{\mathcal{C}}(X)\frac{\operatorname{Tr}\left(U_{\perp}^{\top}H_{X}(V_{s})U_{\perp}\right)}{\operatorname{Tr}\left(\operatorname{Cov}\left(V_{s}^{\top}G(X)\right)\right)}$ $\leq S^{\textup{cl}}(U_{r}|V_{s})\leq 1-\overline{c}(X)\frac{\left\|V_{s}^{\top}\mathbb{E}\left[\nabla G(X)\right]U_{\perp}\right\|_{F}^{2}}{\operatorname{Tr}\left(\operatorname{Cov}\left(V_{s}^{\top}G(X)\right)\right)},$
(29) $\overline{c}(X)\frac{\left\|V_{s}^{\top}\mathbb{E}\left[\nabla G(X)\right]U_{r}\right\|_{F}^{2}}{\operatorname{Tr}\left(\operatorname{Cov}\left(V_{s}^{\top}G(X)\right)\right)}$ $\leq S^{\textup{tot}}(U_{r}|V_{s})\leq\overline{\mathcal{C}}(X)\frac{\operatorname{Tr}\left(U_{r}^{\top}H_{X}(V_{s})U_{r}\right)}{\operatorname{Tr}\left(\operatorname{Cov}\left(V_{s}^{\top}G(X)\right)\right)},$
where $U_{\perp}\in\mathbb{R}^{d\times(d-r)}$ is the orthogonal complement to $U_{r}$.
@@
Our method generalises the derivative-based global sensitivity measures (DGSM) introduced in [50] which are defined as
@@
for a scalar function $f\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$. Setting $V_{s}\in\mathbb{R}^{m\times 1}$ to be a vector, the diagonal elements $H_{X}(V_{s})_{ii}$ correspond to the DGSM $\nu_{i}$ of the scalar function $V_{s}^{\top}G(X)$. Our approach additionally enables sensitivity measures for goal-oriented vector-valued outputs and is not restricted to single-element sets $\tau=\{i\}$. The authors in [35] also derived the following DGSM-based bounds for the total Sobol’ index where $X_{i}\sim\mathcal{N}(\mu_{i},\sigma_{i})$

### v2

@@ Abstract
We introduce a new method to jointly reduce the dimension of the input and output space of a function between high-dimensional spaces. Choosing a reduced input subspace influences which output subspace is relevant and vice versa. Conventional methods focus on reducing either the input or output space, even though both are often reduced simultaneously in practice. Our coupled approach naturally supports goal-oriented dimension reduction, where either an input or output quantity of interest is prescribed. We consider, in particular, goal-oriented sensor placement and goal-oriented sensitivity analysis, which can be viewed as dimension reduction where the most important output or, respectively, input components are chosen. Both applications present difficult combinatorial optimization problems with expensive objectives such as the expected information gain and Sobol’ indices. By optimizing gradient-based bounds, we can determine the most informative sensors and most influential parameters as the largest diagonal entries of some diagnostic matrices, thus bypassing the combinatorial optimization and objective evaluation.
@@ 2 Problem Setting and Objectives
with $\operatorname{\mathbb{D}_{KL}}(\nu\|\mu)=\int\log(\mathop{}\!\mathrm{d}\nu/\mathop{}\!\mathrm{d}\mu)\mathop{}\!\mathrm{d}\nu$. For (O2) to be close to 00, the parameter $X$ needs to be essentially independent of the observations in the subspace $\operatorname{Ker}(V_{s})$, and observations of $Y$ should primarily provide information within the parameter subspace $\operatorname{Im}(U_{r})$. To sample from the approximate density $\widetilde{\pi}_{X|Y}$, we can independently sample $X_{r}\in\mathbb{R}^{r}$, which depends on the projected observation (see algorithms to sample the reduced posterior in [13, Section 6]), and its orthogonal complement $X_{\perp}\in\mathbb{R}^{d-r}$, which only depends on the prior.
Goal-oriented Bayesian optimal experimental design
Bayesian optimal experimental design (BOED) aims to select a set of sensors $\tau\subseteq\{1,\ldots,m\}$ with $|\tau|=s$ that maximizes the information gained about the parameter that is to be inferred. Measurements from the sensors $\tau$ are gathered in the vector $Y_{\tau}=V_{\tau}^{\top}Y$, where $V_{\tau}=[e_{\tau}^{m}]\in\mathbb{R}^{m\times s}$ contains the canonical basis vectors $e_{i}^{m}\in\mathbb{R}^{m}$ indexed by $\tau$. The information content of the experimental design $V_{\tau}$ can be quantified by the expected information gain (EIG), which is defined as the KL-divergence from the prior to the posterior in expectation over the data $\Phi(V_{\tau})=\mathbb{E}_{Y}[\operatorname{\mathbb{D}_{KL}}(\pi_{X|Y_{\tau}}\|\,\pi_{X})]$. When the inference problem focuses on a specific linear quantity of interest in the parameter $X_{r}=U_{r}^{\top}X$, the sensor placement can be improved by maximizing the goal-oriented EIG
(O3) $\max_{V_{\tau}}\;\Phi(V_{\tau}|U_{r})=\max_{V_{\tau}}\;\mathbb{E}_{Y}\left[\operatorname{\mathbb{D}_{KL}}\left(\pi_{X_{r}|Y_{\tau}}\|\,\pi_{X_{r}}\right)\right],$
@@
Goal-oriented global sensitivity analysis
Assuming independent inputs $X_{i}$, $i=1,\ldots,d$, global sensitivity analysis (GSA) aims to identify which set $\tau\subseteq\{1,\ldots,d\}$ of $r=|\tau|$ input factors causes the least output variations in order to fix them to a nominal value and simplify the model. A widely used sensitivity index is the total Sobol’ index $S_{\tau}^{\textup{tot}}=1-\frac{\operatorname{Tr}(\operatorname{Cov}(G(X)|X_{-\tau}))}{\operatorname{Tr}(\operatorname{Cov}(G(X)))}$ with $X_{-\tau}={U_{-\tau}}^{\top}X\in\mathbb{R}^{d-r}$ where $U_{-\tau}=[e_{-\tau}^{d}]\in\mathbb{R}^{d\times(d-r)}$ contains the canonical basis vectors $e_{i}^{d}\in\mathbb{R}^{d}$ indexed by the complement of $\tau$ [28]. A smaller $S_{\tau}^{\textup{tot}}$ implies that the output is less sensitive to the parameters indexed by $\tau$. Thus, the optimal set of input factors $\tau$ for factor fixing minimize the total Sobol’ index $S_{\tau}^{\textup{tot}}$. When there are particular output directions of interest $V_{s}^{\top}G(X)$, it is more suitable to consider minimizing the goal-oriented total Sobol’ index
(O4) $\min_{U_{\tau}}\;S^{\textup{tot}}(U_{\tau}|V_{s})=\min_{U_{\tau}}\;1-\frac{\operatorname{Tr}\left(\operatorname{Cov}\left(\left.V_{s}^{\top}G(X)\right|U_{-\tau}^{\top}X\right)\right)}{\operatorname{Tr}\left(\operatorname{Cov}\left(V_{s}^{\top}G(X)\right)\right)},$
@@ 3 Coupled Input-Output Dimension Reduction
Proposition 3.3 allows us to obtain optimal $U_{r}$ and $V_{s}$ by minimizing the $L^{2}$-error of the forward model as given by (5), the minimum value for the objective in (O1). After that, one can approximate the low-dimensional function $g^{*}$ using state-of-the-art approximation methods (see e.g. [2, 30]) and compute $G^{*}_{\perp}$ to obtain a reduced order model of $G$. We focus on reducing the inputs and/or outputs by minimizing (5) over $U_{r}$ and/or $V_{s}$, which is equivalent to
@@ 3.1 Derivative-based Bounds
We will show in the next subsection that maximizing over either $U_{r}$ or $V_{s}$ yields closed-form solutions that can be computed as dominant eigenvectors of certain diagnostic matrices. There is, however, no closed-form solution for the joint maximization over $U_{r}$ and $V_{s}$. Nonetheless, since the Stiefel manifold $\text{St}(p,q)=\{A\in\mathbb{R}^{p\times q}\;|\;A^{\top}A=I_{q}\}$ is compact [1] and the objective function is continuous with respect to $U_{r}$ and $V_{s}$, the joint maximization problem has a solution. We propose a natural heuristic for its computation in the next subsection.
The following proposition gives sufficient conditions on the density $\pi_{X}$ which ensures $\overline{c}(X)>0$ and $\overline{\mathcal{C}}(X)<\infty$.
@@
While assumption (13) is a classical strict log-concavity assumption on $\pi_{X}$, the assumption (14) means that the density $\pi_{X}$ has to lie above some scaled Gaussian density (see Remark D.5). Assuming access to $\overline{c}(X)$, the lower-bound (11) can be evaluated with no additional cost after computing $U_{r}$ and/or $V_{s}$ and allows us to certify the error made by optimizing the upper bound. For Gaussian priors, $\overline{c}(X)$ can be shown to be the smallest eigenvalue of the prior covariance matrix. For non-Gaussian priors, one may normalize the prior and consider an inference problem in a transformed coordinate system in order to obtain the values for both constants $\overline{\mathcal{C}}(X)$ and $\overline{c}(X)$ [22].
@@ 3.2 Minimizing the Error Bound
Computing these matrix-vector products requires only evaluating the linear-tangent operator $u\mapsto\nabla G(X^{(i)})u$ and its adjoint $v\mapsto\nabla G(X^{(i)})^{\top}v$, which have a computational complexity of roughly two forward simulations [49, 31]. Since the gradient-based dimension reduction methods are intended as a preprocessing step for the analysis of $G$ (e.g., optimization, calibration or parameter estimation), access to gradients is often required by the prospective analysis itself. Finally, if closed-form gradients are not available, automatic differentiation or gradient approximation methods, such as finite differences, may be used. Both the accuracy of the gradient evaluations and the Monte Carlo estimation of the diagnostic matrices will affect the obtained subspaces $U_{r}$ and $V_{s}$. A detailed error analysis is left for future work.
@@
In a recent paper [13], the authors proposed a joint dimension reduction method for the parameter and data space based on similar diagnostic matrices. Using an information theoretic approach, they derive an upper bound for the posterior error $\mathbb{E}_{Y}[\operatorname{\mathbb{D}_{KL}}(\pi_{X|Y}\|\widetilde{\pi}_{X|Y})]$ which is minimized by $U_{r}^{*}$ and $V_{s}^{*}$ consisting of the dominant eigenvectors of
@@ 4 Application to Goal-oriented Coordinate Selection
4.1 Goal-Oriented Bayesian Optimal Experimental Design
@@
Estimating the EIG poses an even greater challenge than solving the Bayesian inference problem as it requires solving the inverse problem itself as a sub-problem. The EIG consists of an expectation over both the posterior and data distributions, necessitating an expensive nested Monte Carlo estimator [55, 15]. Efficient EIG estimation is an active research area, and various approaches exist. Some rely on constructing tractable approximations to the posterior distribution using measure transport [38, 12, 42] or variational inference [26]. Others employ different types of surrogate models [36, 60, 8]. A widely used strategy is to resort to Gaussian linear models for which a closed-form expression of the EIG exists [4, 59, 10]. In addition to the challenging EIG estimation, identifying the optimal set of sensor placements requires solving a combinatorial optimization problem with $\binom{m}{s}$ possible combinations. This NP-hard problem is intractable even for moderate dimensions. Hence, sub-optimal solutions are commonly identified using greedy methods [40, 17, 59] or by solving continuous relaxations of the problem with sparsity constraints [32, 10].
We consider, in particular, goal-oriented BOEDs, where the final target is to infer the parameter in a lower-dimensional linear subspace of interest. This quantity of interest $X_{r}=U_{r}^{\top}X$ is characterized by a prescribed linear ’goal’ operator $U_{r}\in\mathbb{R}^{d\times r}$. For example, if the PDE model describes a spatio-temporal field with $X$ as the initial state, the initial state in a localized region of the domain could be a parameter of interest. Tailoring the experimental design to maximize the information gained with respect to the parameter of interest allows for more efficient and accurate inference from experimental data. For a given parameter of interest $X_{r}$, we seek the optimal design $V_{\star}$ that maximizes the goal-oriented EIG (see (O3))
@@
(24) $\max_{\begin{subarray}{c}V_{\tau}=[e_{\tau}]\in\mathbb{R}^{m\times s}\\ \tau\subset\{1,\ldots,m\},\;|\tau|=s\end{subarray}}\;\operatorname{Tr}\left(V_{\tau}^{\top}H_{Y}(U_{r})V_{\tau}\right)=\sum_{i\in\tau^{*}}H_{Y}(U_{r})_{ii},\textasciiacute$
where the solution $\tau^{*}\subseteq\{1,\ldots,m\}$ is the index set of the $s$ largest diagonal entries of $H_{Y}(U_{r})$. While our solution $\tau^{*}$ may be sub-optimal due to the maximization of a lower bound, it has the advantage of being computable without the expensive evaluation of the EIG nor the need to solve a combinatorial optimization problem.
@@
The EIM takes as input a set of basis vectors $v_{1},\ldots,v_{s}\in\mathbb{R}^{m}$ and returns a set of indices $\{i_{1},\ldots,i_{s}\}\subseteq\{1,\ldots,m\}$ corresponding to the empirically-best interpolation points for the given basis. Combining the EIM with dimension reduction methods for sensor placement has been considered, for example, in [29, 9]. By using the EIM on the columns of $V_{s}^{*}(U_{r})$, we can thus associate sensor locations $\widetilde{\tau}=\{i_{1},\ldots,i_{s}\}$ to the orthogonal design matrix $V_{s}^{*}(U_{r})$. Compared to $\tau^{*}$, the EIM-based sensors $\widetilde{\tau}$ are less optimal with respect to the goal-oriented EIG lower bound (24). However, our numerical experiments show that the EIM-based sensors represent a viable alternative with similarly good performance as $\tau^{*}$ in terms of the true goal-oriented EIG rather than the lower bound. In contrast to $\tau^{*}$, the EIM-based sensors $\widetilde{\tau}$ tend to be less localized in our experiments. In the numerical experiments, we also apply the EIM on PCA bases to obtain sensor placements $\widetilde{\tau}^{\textup{PCA}}$, which we will refer to as PCA-EIM sensors.
4.2 Goal-Oriented Global Sensitivity Analysis
Global sensitivity analysis aims to assign an importance value to each set of input variables, reflecting its contribution to the output variations. It is often used for factor prioritization and/or fixing in complex models. Factor prioritization identifies the subset $\tau\subseteq\{1,\ldots,d\}$, $|\tau|=r$ of input parameters causing the most output variation so that the parameters $X_{\tau}=(X_{i})_{i\in\tau}$ can be prioritized e.g., in uncertainty reduction. Factor fixing, on the other hand, identifies the input set $\tau$ with the least influence on the output. Thus, $X_{\tau}$ can be fixed at a nominal value to reduce the complexity of, e.g., parameter estimation problems. For a scalar-valued square-integrable function $f\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$ with independent inputs $X_{1},\ldots,X_{d}$, commonly used sensitivity indices are the closed Sobol’ indices and the total Sobol’ indices defined as
@@ 4.2 Goal-Oriented Global Sensitivity Analysis
respectively, where $X_{-\tau}=(X_{i})_{i\notin\tau}$. The closed Sobol’ index quantifies the portion of output variance attributed to $X_{\tau}$. A higher value implies a stronger influence of $X_{\tau}$ on the output. The total Sobol’ index encompasses output variation due to $X_{\tau}$ along with variations caused by interactions of $X_{\tau}$ with any of the remaining input variables. As a result, it holds that $S^{\textup{cl}}_{\tau}\leq S^{\textup{tot}}_{\tau}$ and $1=S^{\textup{tot}}_{\tau}+S^{\textup{cl}}_{-\tau}$. A lower value in total Sobol’ index suggests that the parameters indexed by $\tau$ have little direct nor indirect impact on the output.
@@
For $V_{s}\in\mathbb{R}^{m\times 1}$ being a vector, the provided indices correspond to the ones for a scalar-valued output (26).
Using the Poincaré and Cramér-Rao inequality, we now derive derivative-based bounds for the goal-oriented Sobol’ indices involving similar derivative-based quantities as in Theorem 3.5.
@@
With the above notations, for any $U_{\tau}=[e_{\tau}^{d}]\in\mathbb{R}^{d\times r}$ and $V_{s}\in\mathbb{R}^{m\times s}$ with orthogonal columns, we have
(28) $1-\overline{\mathcal{C}}(X)\frac{\operatorname{Tr}\left(U_{-\tau}^{\top}H_{X}(V_{s})U_{-\tau}\right)}{\operatorname{Tr}\left(\operatorname{Cov}\left(V_{s}^{\top}G(X)\right)\right)}$ $\leq S^{\textup{cl}}(U_{\tau}|V_{s})\leq 1-\overline{c}(X)\frac{\left\|V_{s}^{\top}\mathbb{E}\left[\nabla G(X)\right]U_{-\tau}\right\|_{F}^{2}}{\operatorname{Tr}\left(\operatorname{Cov}\left(V_{s}^{\top}G(X)\right)\right)},$
(29) $\overline{c}(X)\frac{\left\|V_{s}^{\top}\mathbb{E}\left[\nabla G(X)\right]U_{\tau}\right\|_{F}^{2}}{\operatorname{Tr}\left(\operatorname{Cov}\left(V_{s}^{\top}G(X)\right)\right)}$ $\leq S^{\textup{tot}}(U_{\tau}|V_{s})\leq\overline{\mathcal{C}}(X)\frac{\operatorname{Tr}\left(U_{\tau}^{\top}H_{X}(V_{s})U_{\tau}\right)}{\operatorname{Tr}\left(\operatorname{Cov}\left(V_{s}^{\top}G(X)\right)\right)},$
where $U_{-\tau}=[e_{-\tau}^{d}]\in\mathbb{R}^{d\times(d-r)}$ is the orthogonal complement to $U_{\tau}$.
@@
Our method generalizes the derivative-based global sensitivity measures (DGSM) introduced in [58] which are defined as
@@
for a scalar function $f\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$. The generalization to vector-valued functions has also been considered in [18]. Setting $V_{s}\in\mathbb{R}^{m\times 1}$ to be a vector, the diagonal elements $H_{X}(V_{s})_{ii}$ correspond to the DGSM $\nu_{i}$ of the scalar function $V_{s}^{\top}G(X)$. Our approach additionally enables sensitivity measures for goal-oriented vector-valued outputs and is not restricted to single-element sets $\tau=\{i\}$. The authors in [41] also derived the following DGSM-based bounds for the total Sobol’ index where $X_{i}\sim\mathcal{N}(\mu_{i},\sigma_{i})$